# On the Number of Control Nodes in Boolean Networks with Degree Constraints

### v1

@@ 3 Boolean networks consisting of XOR functions
According to Theorem 2, if $t^{*}=1$, then the size of the control node set is $n$. On the other hand, it does not mean that $t^{*}$ can take any value from $[3,n]$.
@@
where the subscript is given with modulo $n=4m$ ($+1$). Clearly, this BN is a 3-3-XOR-BN. From Lemma 14 of [27], the entropy is preserved from $[x_{4h+1}(t),x_{4h+2}(t),x_{4h+3}(t),x_{4h+4}(t)]$ to $[x_{4h+5}(t+1),x_{4h+6}(t+1),x_{4h+7}(t+1),x_{4h+8}(t+1)]$. We let $U=\{x_{1},x_{2},x_{3},x_{4}\}$. Then, we can drive the BN to any desired state at time step $m=\frac{n}{4}$ by giving appropriate signal to $[u_{1}(t),u_{2}(t),u_{3}(t),u_{4}(t)]$. By generalizing this construction, we have:
@@
Suppose that $n=[k(k-1)+1]m$, where $k\geq 2$ and $(m\mod k)=0$. Then, for any $k$-$k$-XOR-BN, there exists a control node set of size $\left[1-\frac{k-1}{k(k^{2}-k+1)}\right]n$.
Similar to Lemma 1, we first use the greedy procedure to find a required $V_{1}$, but the difference is that there are at most $k(k-1)$ such $x_{j}$s at STEP 4 because of the constraints on indegree and outdegree, so the size of $S$ decreases by at most $k(k-1)+1$ for each iteration. Hence, we can get a required $V_{1}\subseteq V$ with $|V_{1}|=\frac{1}{k(k-1)+1}n=m$ such that $\Gamma^{+}(x_{i})\cap\Gamma^{+}(x_{j})=\emptyset$ holds for any $x_{i}\neq x_{j}$ in $V_{1}$. Then, we also partition $V$ into three disjoint sets $V_{1},V_{2}$, and $V_{3}$, where $V_{2}=\{x_{j}\mid(x_{i},x_{j})\in E~{}\land~{}x_{i}\in V_{1}~{}\land~{}x_{j}\notin V_{1}\}$ and $V_{3}=V\setminus(V_{1}\cup V_{2})$. Here, $|V_{1}|=m$, $|V_{2}|\geq(k-1)m$ and $|V_{3}|\leq[k(k-2)+1]m$.
@@ 4 Boolean networks consisting of simple AND/OR functions
We construct a simple 2-2-AND-BN by
$x_{2i-1}(t+1)$ $=$ $x_{2i-1}(t)\land x_{2i}(t+1),$
$x_{2i}(t+1)$ $=$ $x_{2i-1}(t)\land x_{2i}(t+1).$
@@
Note that for any node $x_{i}$ in a simple 2-2-AND-BN, we have $|\Gamma^{+}(x_{i})|=|\Gamma^{-}(x_{i})|=2$.
Then for any $x_{i}\notin U$, there are five cases:
$|\Gamma^{+}(x_{i})\cap U|=2$;
$|\Gamma^{+}(x_{i})\cap U|=1$ and $\Gamma^{+}(x_{i})\cap\{x_{i}\}=\emptyset$;
$|\Gamma^{+}(x_{i})\cap U|=1$ and $x_{i}\in\Gamma^{+}(x_{i})$;
$\Gamma^{+}(x_{i})\cap U=\emptyset$ and $\Gamma^{+}(x_{i})\cap\{x_{i}\}=\emptyset$;
$\Gamma^{+}(x_{i})\cap U=\emptyset$ and $x_{i}\in\Gamma^{+}(x_{i})$.
First, it is seen that if a simple 2-2-AND-BN is controllable, then any node $x_{i}\notin U$ will not encounter Cases (iii) and (v), that is, $x_{i}\notin\Gamma^{+}(x_{i})$. Suppose there is a node $x_{i}\notin U$, where $x_{i}\in\Gamma^{+}(x_{i})$. If $x_{i}(0)=0$, then $x_{i}(t)$ will always be 0, which contradicts the fact that the simple 2-2-AND-BN is controllable.
@@
which implies that $|V_{0}|+|V_{1}|+|V_{2}|\leq 2|U|$ and then
$|V|$ $=$ $|V_{0}|+|V_{1}|+|V_{2}|+|U|$
$\leq$ $3|U|.$
Therefore, in order to satisfy the controllability condition, we need to control at least $\frac{n}{3}$ nodes. $\Box$
@@
According to Eqs. (1) and (2), we have
$2|U|$ $=$ $(|V_{1}|+|W_{1}|)+2(|V_{2}|+|W_{2}|)$
$\geq$ $(n-|U|)+2(|V_{2}|+|W_{2}|)\quad({\rm By}~{}|V_{0}|+|V_{1}|+|V_{2}|=n-|U|),$
and then $|U|\geq\frac{1}{3}n+\frac{2}{3}(|V_{2}|+|W_{2}|)$. It can be seen that $|U|=\frac{1}{3}n$ only when $|V_{2}|+|W_{2}|=0$ and $|V_{1}|+|W_{1}|=n-|U|=\frac{2}{3}n$.
Furthermore, based on $|U|=|W_{0}|+|W_{1}|+|W_{2}|$ and $|V|=|V_{0}|+|V_{1}|+|V_{2}|+|U|$, we deduce that $|U|=\frac{1}{3}n$ only when $|V_{0}|+|W_{0}|=\frac{1}{3}n,~{}|V_{1}|+|W_{1}|=\frac{2}{3}n,~{}|V_{2}|+|W_{2}|=0$ and $|V_{0}|=|W_{1}|$.
@@
For any integer $n=(2k-1)m+1$ with a positive integer $m>1$, there exists a simple $k$-$k$-AND-BN for which the size of the minimum control node set is $(k-1)m+1=\frac{k-1}{2k-1}(n-1)+1$.
We construct a simple $k$-$k$-AND-BN by
@@
$x_{(2k-1)m-k+3}(t+1)$ $=$ $x_{km+1}(t)\land x_{km+2}(t)\land\cdots\land x_{km+m-2}(t)\land x_{km+m-1}(t)\land x_{km+m}(t),$
$x_{(2k-1)m-k+4}(t+1)$ $=$ $x_{km+1}(t)\land x_{km+2}(t)\land\cdots\land x_{km+m-2}(t)\land x_{km+m-1}(t)\land x_{km+m}(t),$
@@
$x_{n}(t+1)$ $=$ $x_{km+1}(t)\land x_{km+2}(t)\land\cdots\land x_{km+m-2}(t)\land x_{km+m-1}(t)\land x_{km+m}(t)$
@@
Note that for any node $x_{i}$ in a simple $k$-$k$-AND-BN, we have $|\Gamma^{+}(x_{i})|=|\Gamma^{-}(x_{i})|=k$.
@@ 5 Boolean networks consisting of AND/OR functions and negations
Similar to Theorem 5, we have
$|V_{1}|+2|V_{2}|+|W_{1}|+2|W_{2}|$ $=$ $2|U|,$ (5)
$|V_{0}|+|V_{1}|+|V_{2}|$$\leq$ $|V_{1}|+|W_{1}|,$ (6)
where $|V|=|V_{0}|+|V_{1}|+|V_{2}|+|U|$ and $|U|=|W_{0}|+|W_{1}|+|W_{2}|$. Here, Eqs. (5) and (6) correspond to Eqs. (1) and (2) in Theorem 5, respectively.
Then we note that for any node $x_{i}\in M_{1}$, $|\Gamma^{+}(x_{i})\cap U|\geq 1$. Suppose there is a node $x_{i}\in M_{1}$, where $|\Gamma^{+}(x_{i})\cap U|=0$. We assume without loss of generality $x_{i_{1}}(t+1)=x_{i}(t)\land x_{k_{1}}(t)$ and $x_{i_{2}}(t+1)=\overline{x_{i}(t)}\land x_{k_{2}}(t)$. Here, $x_{i_{1}},x_{i_{2}}\notin U$, because $|\Gamma^{+}(x_{i})\cap U|=0$. Then for any initial state ${\bf x}^{0}$, the target state of the form $x_{i_{1}}=x_{i_{2}}=1$ cannot be reached, which contradicts the fact that the 2-2-AND-BN with negation is controllable.
@@
  • When $n-|U|>|M_{1}|$, we can minimize $2|U|=(|V_{1}|+|W_{1}|)+2(|V_{2}|+|W_{2}|)$ to $n-|U|$, where $M_{1}\subseteq V_{1}\cup W_{1},~{}|V_{1}|+|W_{1}|=n-|U|$ and $|V_{2}|+|W_{2}|=0$, while satisfying the above two conditions.
@@
Define $M_{j},~{}j=0,1,\ldots,\lfloor\frac{k}{2}\rfloor$ as the set of nodes where the literal $x_{i}$ or $\overline{x_{i}}$ ($i=1,2,\ldots,n$) appears $j$ times in the $k$-$k$-AND-BN with negation. Since the outdegree of each node in the $k$-$k$-AND-BN with negation is $k$, for any node $x_{i}\in M_{j}$, there are two cases, one is that the literal $x_{i}$ occurs $j$ times and the literal $\overline{x_{i}}$ occurs $k-j$ times; the other is that the literal $x_{i}$ occurs $k-j$ times and the literal $\overline{x_{i}}$ occurs $j$ times. For example, if $x_{1}(t+1)=\overline{x_{1}(t)}\land x_{2}(t)\land\overline{x_{3}(t)},~{}x_{2}(t+1)=x_{1}(t)\land\overline{x_{2}(t)}\land\overline{x_{3}(t)},~{}x_{3}(t+1)=x_{1}(t)\land\overline{x_{2}(t)}\land\overline{x_{3}(t)}$, then $M_{0}=\{x_{3}\}$ and $M_{1}=\{x_{1},x_{2}\}$.
@@
(Proof) Similar to Theorem 6, we have
$|V_{0}|+|V_{1}|+\cdots+|V_{k}|$$\leq$ $|V_{k-1}|+|W_{k-1}|,$ (9)
where $|V|=|V_{0}|+|V_{1}|+\cdots+|V_{k}|+|U|$ and $|U|=|W_{0}|+|W_{1}|+\cdots+|W_{k}|$. Here, Eqs. (8) and (9) correspond to Eqs. (3) and (4) in Theorem 6, respectively.
@@
  • When $n-|U|>\sum_{j=1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|$, we can minimize $k|U|=(|V_{1}|+|W_{1}|)+\cdots+(k-1)(|V_{k-1}|+|W_{k-1}|)+k(|V_{k}|+|W_{k}|)$ to $(k-1)(n-|U|)$, where $\cup_{j=1}^{\lfloor\frac{k}{2}\rfloor}M_{j}\subseteq V_{k-1}\cup W_{k-1},~{}|V_{k-1}|+|W_{k-1}|=n-|U|$ and $|V_{1}|+|W_{1}|=\cdots=|V_{k-2}|+|W_{k-2}|=|V_{k}|+|W_{k}|=0$, while satisfying the above two conditions.
@@
00 0 $\geq(2k-l)\left[\frac{k-l}{2k-l}n+\frac{1}{2k-l}\left(\sum_{j=1}^{l-2}j\cdot|M_{j}|\right)+\frac{l-1}{2k-l}\left(\sum_{j=l-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)-|U|\right]$
@@
It can be seen that $U=\{x_{1},x_{2},x_{3},x_{5},x_{12},x_{13},x_{14},x_{15}\}$ ($|U|=8$) is the minimum control node set. Suppose there is a control node set $U^{\prime}$, where $|U^{\prime}|=7$, then according to the conditions (I) and (II) in the proof of Theorem 8, we need to satisfy $|V_{3}|+|W_{3}|\geq 8$ and for any node $x_{i}\in M_{1}$, $|\Gamma^{+}(x_{i})\cap U|\geq 1$. Hence, the value of $(|V_{1}|+|W_{1}|)+2(|V_{2}|+|W_{2}|)+3(|V_{3}|+|W_{3}|)$ is not less than $(15-8)+3\times 8=31$. However, $314|U^{\prime}|=28<$ 31, which contradicts Eq. (8), so $U=\{x_{1},x_{2},x_{3},x_{5},x_{12},x_{13},x_{14},x_{15}\}$ is the minimum control node set.
@@
It can be seen that $U=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{12},x_{13},x_{14},x_{15}\}$ ($|U|=9$) is the minimum control node set. Suppose there is a control node set $U^{\prime}$, where $|U^{\prime}|=8$, then according to the conditions (I) and (II) in the proof of Theorem 8, we need to satisfy $|V_{3}|+|W_{3}|\geq 7$ and for any node $x_{i}\in M_{1}$ (resp., $x_{i}\in M_{2}$), $|\Gamma^{+}(x_{i})\cap U|\geq 1$ (reps. $|\Gamma^{+}(x_{i})\cap U|\geq 2$). Hence, the value of $(|V_{1}|+|W_{1}|)+2(|V_{2}|+|W_{2}|)+3(|V_{3}|+|W_{3}|)$ is not less than $3+2\times(12-7)+3\times 7=34$. However, $344|U^{\prime}|=32<$ 34, which contradicts Eq. (8), so $U=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{12},x_{13},x_{14},x_{15}\}$ is the minimum control node set.

### v2

@@ 3 Boolean networks consisting of XOR functions
According to Theorem 2, if $t^{*}=1$, then the size of the control node set is $n$. On the other hand, it does not mean that $t^{*}$ can take any value from $[3,n]$. To illustrate why $t^{*}$ cannot take any value from $[3,n]$, we consider the 2-2-XOR-BN in Example 1, where $n=6$ and
$x_{1}(t+1)$ $=$ $x_{1}(t)\oplus x_{3}(t),$
$x_{2}(t+1)$ $=$ $x_{1}(t)\oplus x_{3}(t),$
$x_{3}(t+1)$ $=$ $x_{2}(t)\oplus x_{4}(t),$
$x_{4}(t+1)$ $=$ $x_{2}(t)\oplus x_{5}(t),$
$x_{5}(t+1)$ $=$ $x_{4}(t)\oplus x_{6}(t),$
$x_{6}(t+1)$ $=$ $x_{5}(t)\oplus x_{6}(t).$
If $t^{*}$ could take the value 6, then $|U_{1}|=1$ according to Theorem 2, and there are six cases as follows (each case can be verified by exhaustive examination of states of $x_{i}(t)$ with $t=0,\ldots,6$ for $x_{i}\in U_{1}$):
  • When $U_{1}=\{x_{1}\}$, it can be seen that we cannot drive this BN from any initial state ${\bf x}^{0}$ to any target state ${\bf x}^{T}$ at $t=6$. For example, let ${\bf x}^{0}=[0,0,0,0,0,0]$ and ${\bf x}^{T}=[0,0,0,1,0,0]$.
  • When $U_{1}=\{x_{2}\}$, it can be seen that we cannot drive this BN from any initial state ${\bf x}^{0}$ to any target state ${\bf x}^{T}$ at $t=6$. For example, let ${\bf x}^{0}=[0,0,0,0,0,0]$ and ${\bf x}^{T}=[0,0,0,1,0,0]$.
  • When $U_{1}=\{x_{3}\}$, it can be seen that we cannot drive this BN from any initial state ${\bf x}^{0}$ to any target state ${\bf x}^{T}$ at $t=6$. For example, let ${\bf x}^{0}=[0,0,0,0,0,0]$ and ${\bf x}^{T}=[1,0,0,0,0,0]$.
  • When $U_{1}=\{x_{4}\}$, it can be seen that we cannot drive this BN from any initial state ${\bf x}^{0}$ to any target state ${\bf x}^{T}$ at $t=6$. For example, let ${\bf x}^{0}=[0,0,0,0,0,0]$ and ${\bf x}^{T}=[0,0,1,0,0,0]$.
  • When $U_{1}=\{x_{5}\}$, it can be seen that we cannot drive this BN from any initial state ${\bf x}^{0}$ to any target state ${\bf x}^{T}$ at $t=6$. For example, let ${\bf x}^{0}=[0,0,0,0,0,0]$ and ${\bf x}^{T}=[0,0,0,1,0,0]$.
  • When $U_{1}=\{x_{6}\}$, it can be seen that we cannot drive this BN from any initial state ${\bf x}^{0}$ to any target state ${\bf x}^{T}$ at $t=6$. For example, let ${\bf x}^{0}=[0,0,0,0,0,0]$ and ${\bf x}^{T}=[1,0,0,0,0,0]$.
Therefore, $t^{*}$ cannot take the value 6.
@@
where the subscript is given with modulo $n=4m$ (+1+1). Clearly, this BN is a 3-3-XOR-BN. From Lemma 14 of [31], the entropy is preserved from $x_{4h+1}(t),x_{4h+2}(t),x_{4h+3}(t),x_{4h+4}(t)]$ to $x_{4h+5}(t+1),x_{4h+6}(t+1),x_{4h+7}(t+1),x_{4h+8}(t+1)]$. We let $U=\{x_{1},x_{2},x_{3},x_{4}\}$. Then, we can drive the BN to any desired state at time step $m=\frac{n}{4}$ by giving appropriate signal to $u_{1}(t),u_{2}(t),u_{3}(t),u_{4}(t)]$. By generalizing this construction, we have:
@@
Suppose that n=[k(k1)+1]mn=[k(k-1)+1]m, where $k\geq 2$ and $(m\mod k)=0$. Then, for any $k$-$k$-XOR-BN, there exists a control node set of size $\left[1-\frac{k-1}{k(k^{2}-k+1)}\right]n$.
(Proof) Similar to Lemma 1, we first use the greedy procedure to find a required $V_{1}$, but the difference is that there are at most k(k1)k(k-1) such $x_{j}$s at STEP 4 because of the constraints on indegree and outdegree, so the size of $S$ decreases by at most k(k1)+1k(k-1)+1 for each iteration. Hence, we can get a required $V_{1}\subseteq V$ with $|V_{1}|=\frac{1}{k(k-1)+1}n=m$ such that $\Gamma^{+}(x_{i})\cap\Gamma^{+}(x_{j})=\emptyset$ holds for any $x_{i}\neq x_{j}$ in $V_{1}$. Then, we also partition $V$ into three disjoint sets $V_{1},V_{2}$, and $V_{3}$, where $V_{2}=\{x_{j}\mid(x_{i},x_{j})\in E~\land~x_{i}\in V_{1}~\land~x_{j}\notin V_{1}\}$ and $V_{3}=V\setminus(V_{1}\cup V_{2})$. Here, $|V_{1}|=m$, $|V_{2}|\geq(k-1)m$ and $|V_{3}|\leq[k(k-2)+1]m$.
@@ 4 Boolean networks consisting of simple AND/OR functions
(Proof) We construct a simple 2-2-AND-BN by
$x_{2i-1}(t+1)$ $=$ $x_{2i-1}(t)\land x_{2i}(t),$
$x_{2i}(t+1)$ $=$ $x_{2i-1}(t)\land x_{2i}(t).$
@@
In the following, we show best case upper bounds for 2-AND-BN and 2-2-AND-BN, where the latter will be generalized for $k$-$k$-AND-BN. The basic idea is to use a kind of shift-register, by which we can propagate signals given to the control node (e.g., $x_{1}$) to the other nodes, where some additional nodes are required to satisfy degree constraints.
@@
(Proof) Note that for any node $x_{i}$ in a simple 2-2-AND-BN, we have $|\Gamma^{+}(x_{i})|=|\Gamma^{-}(x_{i})|=2$.
Then for any $x_{i}\notin U$, there are five cases: (i) $|\Gamma^{+}(x_{i})\cap U|=2$; (ii) $|\Gamma^{+}(x_{i})\cap U|=1$ and $\Gamma^{+}(x_{i})\cap\{x_{i}\}=\emptyset$; (iii) $|\Gamma^{+}(x_{i})\cap U|=1$ and $x_{i}\in\Gamma^{+}(x_{i})$; (iv) $\Gamma^{+}(x_{i})\cap U=\emptyset$ and $\Gamma^{+}(x_{i})\cap\{x_{i}\}=\emptyset$; (v) $\Gamma^{+}(x_{i})\cap U=\emptyset$ and $x_{i}\in\Gamma^{+}(x_{i})$. First, it is seen that if a simple 2-2-AND-BN is controllable, then any node $x_{i}\notin U$ will not encounter Cases (iii) and (v), that is, $x_{i}\notin\Gamma^{+}(x_{i})$. Suppose there is a node $x_{i}\notin U$, where $x_{i}\in\Gamma^{+}(x_{i})$. If $x_{i}(0)=0$, then $x_{i}(t)$ will always be 0, which contradicts the fact that the simple 2-2-AND-BN is controllable.
@@
which implies that $|V_{0}|+|V_{1}|+|V_{2}|\leq 2|U|$ and then $|V|=|V_{0}|+|V_{1}|+|V_{2}|+|U|\leq 3|U|$. Therefore, in order to satisfy the controllability condition, we need to control at least $\frac{n}{3}$ nodes. $\Box$
@@
To further illustrate how Eqs. (1) and (2) are derived, we construct the following 2-2-AND-BN (see Fig. 4),
Figure 4: A 2-2-AND-BN, where the nodes in the top represent those at time $t$ and the nodes in the bottom represent those at time $t+1$.
$x_{1}(t+1)$ $=$ $x_{2}(t)\land x_{3}(t),$
$x_{2}(t+1)$ $=$ $x_{4}(t)\land x_{8}(t),$
$x_{3}(t+1)$ $=$ $x_{1}(t)\land x_{9}(t),$
$x_{4}(t+1)$ $=$ $x_{1}(t)\land x_{10}(t),$
$x_{5}(t+1)$ $=$ $x_{7}(t)\land x_{11}(t),$
$x_{6}(t+1)$ $=$ $x_{7}(t)\land x_{12}(t),$
$x_{7}(t+1)$ $=$ $x_{2}(t)\land x_{5}(t),$
$x_{8}(t+1)$ $=$ $x_{3}(t)\land x_{5}(t),$
$x_{9}(t+1)$ $=$ $x_{4}(t)\land x_{6}(t),$
$x_{10}(t+1)$ $=$ $x_{6}(t)\land x_{8}(t),$
$x_{11}(t+1)$ $=$ $x_{9}(t)\land x_{13}(t),$
$x_{12}(t+1)$ $=$ $x_{10}(t)\land x_{13}(t),$
$x_{13}(t+1)$ $=$ $x_{11}(t)\land x_{12}(t),$
where $V_{0}=\{x_{1}\}$, $V_{1}=\{x_{2},x_{3},x_{4}\}$, $V_{2}=\{x_{5},x_{6}\}$, $W_{0}=\{x_{7}\}$, $W_{1}=\{x_{8},x_{9},x_{10},x_{11},x_{12}\}$, and $W_{2}=\{x_{13}\}$.
It is clear that $V_{0},V_{1},V_{2},W_{0},W_{1},W_{2}$ satisfy the conditions stated in Theorem 5, where $V_{j}=\{x_{i}\mid x_{i}\notin U~\land~|\Gamma^{+}(x_{i})\cap U|=j~\land~\Gamma^{+}(x_{i})\cap\{x_{i}\}=\emptyset~\land~\Gamma^{-}(x_{i})\cap(V_{1}\cup W_{1})\neq\emptyset\}$, $W_{j}=\{x_{i}\mid x_{i}\in U~\land~|\Gamma^{+}(x_{i})\cap U|=j\}$, $j=0,1,2$, and $U=W_{0}\cup W_{1}\cup W_{2}$.
Then, for Eq. (1), we can see the corresponding result in Fig. 5 because each node in $V_{1}\cup W_{1}$ has one edge (green bold line) to $U$, each node in $V_{2}\cup W_{2}$ has two edges (red bold lines) to $U$, and $U$ have edges only from $V_{1}\cup V_{2}\cup W_{1}\cup W_{2}$.
Figure 5: Illustration of Eq. (1).
In addition, for Eq. (2), we can see the corresponding result in Fig. 6 because each node in $V_{0}\cup V_{1}\cup V_{2}$ has at least one edge (blue bold line) from $V_{1}\cup W_{1}$ (note that $\Gamma^{-}(x_{i})\cap(V_{1}\cup W_{1})\neq\emptyset$ holds for any node $x_{i}\notin U$).
Figure 6: Illustration of Eq. (2).
@@
For any integer n=(2k1)m+1n=(2k-1)m+1 with a positive integer $m>1$, there exists a simple $k$-$k$-AND-BN for which the size of the minimum control node set is $(k-1)m+1=\frac{k-1}{2k-1}(n-1)+1$.
(Proof) We construct a simple $k$-$k$-AND-BN by
@@
$x_{(2k-1)m-k+3}(t+1)$ $=$ $x_{km+1}(t)\land x_{km+2}(t)\land\cdots\land x_{km+k-2}(t)\land x_{km+k-1}(t)\land x_{km+k}(t),$
$x_{(2k-1)m-k+4}(t+1)$ $=$ $x_{km+1}(t)\land x_{km+2}(t)\land\cdots\land x_{km+k-2}(t)\land x_{km+k-1}(t)\land x_{km+k}(t),$
@@
$x_{n}(t+1)$ $=$ $x_{km+1}(t)\land x_{km+2}(t)\land\cdots\land x_{km+k-2}(t)\land x_{km+k-1}(t)\land x_{km+k}(t)$
@@
(Proof) Note that for any node $x_{i}$ in a simple $k$-$k$-AND-BN, we have $|\Gamma^{+}(x_{i})|=|\Gamma^{-}(x_{i})|=k$.
@@ 5 Boolean networks consisting of AND/OR functions and negations
(Proof) First, the relations given by Eqs. (1) and (2) in Theorem 5 remain valid in this case. Then we note that for any node $x_{i}\in M_{1}$, $|\Gamma^{+}(x_{i})\cap U|\geq 1$. Suppose there is a node $x_{i}\in M_{1}$, where $|\Gamma^{+}(x_{i})\cap U|=0$. We assume without loss of generality $x_{i_{1}}(t+1)=x_{i}(t)\land x_{k_{1}}(t)$ and $x_{i_{2}}(t+1)=\overline{x_{i}(t)}\land x_{k_{2}}(t)$. Here, $x_{i_{1}},x_{i_{2}}\notin U$, because $|\Gamma^{+}(x_{i})\cap U|=0$. Then for any initial state ${\bf x}^{0}$, the target state of the form $x_{i_{1}}=x_{i_{2}}=1$ cannot be reached, which contradicts the fact that the 2-2-AND-BN with negation is controllable.
@@
  • When $n-|U|>|M_{1}|$, we can minimize $2|U|=(|V_{1}|+|W_{1}|)+2(|V_{2}|+|W_{2}|)$ to n|U|n-|U|, where $M_{1}\subseteq V_{1}\cup W_{1},~|V_{1}|+|W_{1}|=n-|U|$ and $|V_{2}|+|W_{2}|=0$, while satisfying the above two conditions.
@@
Define $M_{j},~j=0,1,\ldots,\lfloor\frac{k}{2}\rfloor$ as the set of nodes where the literal $x_{i}$ or $\overline{x_{i}}$ ($i=1,2,\ldots,n$) appears $j$ times in the $k$-$k$-AND-BN with negation. Since the outdegree of each node in the $k$-$k$-AND-BN with negation is $k$, for any node $x_{i}\in M_{j}$, there are two cases, one is that the literal $x_{i}$ occurs $j$ times and the literal $\overline{x_{i}}$ occurs kjk-j times; the other is that the literal $x_{i}$ occurs kjk-j times and the literal $\overline{x_{i}}$ occurs $j$ times. For example, if $x_{1}(t+1)=\overline{x_{1}(t)}\land x_{2}(t)\land\overline{x_{3}(t)},~x_{2}(t+1)=x_{1}(t)\land\overline{x_{2}(t)}\land\overline{x_{3}(t)},~x_{3}(t+1)=x_{1}(t)\land\overline{x_{2}(t)}\land\overline{x_{3}(t)}$, then $M_{0}=\{x_{3}\}$ and $M_{1}=\{x_{1},x_{2}\}$.
@@
(Proof) First, the relations given by Eqs. (3) and (4) in Theorem 6 remain valid in this case.
@@
  • When $n-|U|>\sum_{j=1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|$, we can minimize $k|U|=(|V_{1}|+|W_{1}|)+\cdots+(k-1)(|V_{k-1}|+|W_{k-1}|)+k(|V_{k}|+|W_{k}|)$ to (k1)(n|U|)(k-1)(n-|U|), where $\cup_{j=1}^{\lfloor\frac{k}{2}\rfloor}M_{j}\subseteq V_{k-1}\cup W_{k-1},~|V_{k-1}|+|W_{k-1}|=n-|U|$ and $|V_{1}|+|W_{1}|=\cdots=|V_{k-2}|+|W_{k-2}|=|V_{k}|+|W_{k}|=0$, while satisfying the above two conditions.
@@
$0$ $\geq(2k-l)\left[\frac{k-l}{2k-l}n+\frac{1}{2k-l}\left(\sum_{j=1}^{l-2}j\cdot|M_{j}|\right)+\frac{l-1}{2k-l}\left(\sum_{j=l-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right)-|U|\right]$
@@
It can be seen that $U=\{x_{1},x_{2},x_{3},x_{5},x_{12},x_{13},x_{14},x_{15}\}$ ($|U|=8$) is the minimum control node set. Suppose there is a control node set $U^{\prime}$, where $|U^{\prime}|=7$, then according to the conditions (I) and (II) in the proof of Theorem 8, we need to satisfy $|V_{3}|+|W_{3}|\geq 8$ and for any node $x_{i}\in M_{1}$, $|\Gamma^{+}(x_{i})\cap U|\geq 1$. Hence, the value of $(|V_{1}|+|W_{1}|)+2(|V_{2}|+|W_{2}|)+3(|V_{3}|+|W_{3}|)$ is not less than $(15-8)+3\times 8=31$. However, $4|U^{\prime}|=28<31$, which contradicts Eq. (3), so $U=\{x_{1},x_{2},x_{3},x_{5},x_{12},x_{13},x_{14},x_{15}\}$ is the minimum control node set.
@@
It can be seen that $U=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{12},x_{13},x_{14},x_{15}\}$ ($|U|=9$) is the minimum control node set. Suppose there is a control node set $U^{\prime}$, where $|U^{\prime}|=8$, then according to the conditions (I) and (II) in the proof of Theorem 8, we need to satisfy $|V_{3}|+|W_{3}|\geq 7$ and for any node $x_{i}\in M_{1}$ (resp., $x_{i}\in M_{2}$), $|\Gamma^{+}(x_{i})\cap U|\geq 1$ (reps. $|\Gamma^{+}(x_{i})\cap U|\geq 2$). Hence, the value of $(|V_{1}|+|W_{1}|)+2(|V_{2}|+|W_{2}|)+3(|V_{3}|+|W_{3}|)$ is not less than $3+2\times(12-7)+3\times 7=34$. However, $4|U^{\prime}|=32<34$, which contradicts Eq. (3), so $U=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{12},x_{13},x_{14},x_{15}\}$ is the minimum control node set.